if u have several 2 micro faraday capacitors each capable of withstanding 200V without breakdown how would you assemble a combination that has an equivalent capacitance pf 0.4 micro faraday withstanding 1000V ?

Grade:12

1 Answers

Arun
4 years ago
Put five such capacitors in series. Then, the equivalent capacitance is 
2.0X10^(-6)F/5 = 0.40X10^(-6)F. 

With each capacitor taking a 200-V potential difference, the equivalent capacitor can withstand 1000 V.  

(b) As one possibility, you can take three identical arrays of capacitors, each array being a five-capacitor combination described in part (a) above, and hook up the arrays in parallel. The equivalent capacitance is now Ceq = 3(0.40X10^(-6)F) = 1.2X10^(-6)F. With each capacitor taking a 200-V potential difference, the equivalent capacitor can withstand 1000 V.